I got this wrong..I'm totally lost thank you

Practice 8.1 example 26:

Find the values for m and b in the following system so that the solution to the system is

(-3, 4)

5x +7y=b
mx+ y =22

Answer by tutmath(14)

In this you need to substitute (-3, 4) in place of (x,y)
In the first equation:
5x +7y=b
5(-3) + 7(4) = b
-15 + 28 = b
b = 13
In the second equation:
mx+ y =22
m(-3) + 4 = 22
-3m = 18
m = -6
(If you have any questions on the solution you can send me your queries at tutmath@gmail.com)